### v1

# Run LoRA Run: Faster and Lighter LoRA Implementations

Abstract
LoRA is a technique that reduces the number of trainable parameters in a neural network by introducing low-rank adapters to linear layers. This technique is used both for fine-tuning (LoRA, QLoRA) and full train (ReLoRA). This paper presents the RunLoRA framework for efficient implementations of LoRA that significantly improves the speed of neural network training and fine-tuning using low-rank adapters. The proposed implementation optimizes the computation of LoRA operations based on dimensions of corresponding linear layer, layer input dimensions and lora rank by choosing best forward and backward computation graph based on FLOPs and time estimations, resulting in faster training without sacrificing accuracy. The experimental results show up to 17% speedup on Llama family of models.
Keywords Low-rank  $\cdot$ LoRA  $\cdot$ Llama
1 Introduction
LoRA Hu et al. (2022) paper has introduced low-rank adapters to fine-tune large LLMs on downstream tasks. This approach quickly became popular due to reduced cost of the update. Different modifications of LoRA followed, for example, QLoRA Dettmers et al. (2023) utilizes quantization and further reduces fine-tuning costs, and ReLoRA Lialin et al. (2023) which showed that low-rank updates can also be used for full training.
However, all variations of LoRA use the same chain of operations while calculating the output, which often leads to sub-optimal graph of computations. We propose RunLora: a framework which contains different variations of forward and backward pass through an adapter-induced linear layer and chooses the best pair for a given architecture.
We evaluated our framework’s performance on a series of Llama models and achieved up to 17% speedup only due to optimized chain of PyTorch operations. Additionally, we managed to save up to 4Gb of memory due to reduction in number of saved activations.
2 Problem setting and Methodology
Default forward pass of LoRA looks the following:
$LoRA(X)=XW+(XA)B,$ (1)
The backward pass is then automatically determined by the framework from an autograd technique. All the optimizations are left to the neural network training framework, which often performs sub-optimally.
Many scientists and engineers avoid the following chain of computations:
$LoRA(X)=X(W+AB).$ (2)
This is due to an implicit assumption that weights $W$ are quite large and forming a same-size matrix AB is undesirable. However, real-world LoRA-adapter training deals with large input $X$ in an attempt of increasing batch size to utilize GPU RAM at its full capacity. Large batch size leads to a contradiction to the assumption and inefficient LoRA implementation.
Our current implementation contains 2 major and 2 minor variants of the forward pass and 5 major and 30 minor variants of the backward pass. Major variants are different by opening brackets differently, while minor variants are based on major variants with operations reordered. Formally, the major forward variants are:
  1. 1.
    $Y=(XW)+((XA)B)$
  2. 2.
    $Y=X(W+AB)$
Unlike default LoRA implementation, neither forward function in RunLoRA saves result of $XA$ to context.
The backward of the LoRA adapter requires us to calculate the following tensors:
$\left\{\begin{array}[]{l}\mathrm{d}A=X^{\top}\mathrm{d}YB^{\top},\\ \mathrm{d}B=A^{\top}X^{\top}\mathrm{d}Y,\\ \mathrm{d}X=\mathrm{d}YW^{\top}+\mathrm{d}YB^{\top}A^{\top}.\end{array}\right.$ (3)
Due to associativity of matrix multiplications, there are several ways of doing it and getting the same result up to rounding errors:
$ABC=(AB)C=A(BC).$ (4)
There are 3 multiplications and each can be done in 2 ways, which leads to 8 variants of the backward pass:
  1. 1.
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
  2. 2.
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
  3. 3.
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
  4. 4.
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
  5. 5.
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
  6. 6.
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
  7. 7.
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
  8. 8.
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
However, out of 8 variants we implement only 5 as we assume others to require more or equal number of FLOPs in any setting. For example, backward6 would require more flops than backward5 for any architecture and training configuration (Table 1).
  1. 1.
    backward1, equivalent to the variant 1 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=XA,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=Z_{2}^{\top}\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  2. 2.
    backward2, equivalent to the variant 2 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  3. 3.
    backward3, equivalent to the variant 3 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=Z_{2}B^{\top},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  4. 4.
    backward4, equivalent to the variant 4 above:
    $Z_{1}=W+AB,Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=Z_{2}B^{\top},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YZ_{1}^{\top},$
  5. 5.
    backward5, equivalent to the variant 5 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=XA,Z_{3}=W+AB,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=Z_{2}^{\top}\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YZ_{3}^{\top}.$
Method FLOPs
forward1 $2b\cdot s\cdot(i\cdot o+r\cdot i+o\cdot r)$
forward2 $2(i\cdot o\cdot r+b\cdot s\cdot o\cdot i)$
forward3 $2b\cdot s\cdot(i\cdot o+r\cdot i+o\cdot r)$
forward4 $2(i\cdot o\cdot r+b\cdot s\cdot o\cdot i)$
backward1 $2b\cdot s\cdot(2o\cdot r+3i\cdot r+o\cdot i)$
backward2 $2b\cdot s(o\cdot r+2i\cdot r+2i\cdot o)+2i\cdot o\cdot r$
backward3 $2b\cdot s\cdot(2i\cdot o+o\cdot r+i\cdot r)+4i\cdot r\cdot o$
backward4 $2(2b\cdot s\cdot i\cdot o+3i\cdot o\cdot r)$
backward5 $2b\cdot s\cdot(2o\cdot r+2i\cdot r+o\cdot i)+2i\cdot o\cdot r$
backward6 $2b\cdot s\cdot(2o\cdot r+2i\cdot r+2o\cdot i)+4i\cdot o\cdot r$
backward7 $2b\cdot s\cdot(o\cdot r+i\cdot r+2o\cdot i)+4i\cdot o\cdot r$
backward8 $2b\cdot s\cdot(o\cdot r+i\cdot r+2o\cdot i)+4i\cdot o\cdot r$
Table 1: Number of floating-point operations per second(FLOPs) for our implemented forward and backward passes. $b$ batch size; $s$ sequence length; $i$ input dimension, $o$ output dimension; $r$ adapter rank;
3 Numerical experiments
We have conducted a number of experiments on LLama family of models ranging from 60 millions of parameters up to 1.3 billion. We measured mean time of a Forward-Backward pass through the network for different architectures and training settings. As showed in Table 3, we managed to achieve up to 17% speedup compared to default peft implementation.
Additionally, since RunLoRA forward functions do not save intermediate result $XA$ we manage to save up to 4Gb of memory.
All experiments were performed on a single Nvidia A100 GPU 80Gb. Training was conducted in bfloat16 format.
Implementation Mean F-B loop, ms Memory for F-B loop, MB Speedup, % Memory Saved, MB
[…]
Table 2: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank.
4 Conclusion and Future Work
One of the possible directions for future work is utilizing quantization of weights to conduct experiments on larger batch sizes and evaluate RunLoRA on some of the current top LLMs, for example, Llama2-7b/13b/70b. We expect larger batch sizes and architectures to show even greater performance improvement.